# Ensemble Kalman Diffusion Guidance: A Derivative-free Method for Inverse Problems

### v1

@@ Abstract
When solving inverse problems, it is increasingly popular to use pre-trained diffusion models as plug-and-play priors. This framework can accommodate different forward models without re-training while preserving the generative capability of diffusion models. Despite their success in many imaging inverse problems, most existing methods rely on privileged information such as derivative, pseudo-inverse, or full knowledge about the forward model. This reliance poses a substantial limitation that restricts their use in a wide range of problems where such information is unavailable, such as in many scientific applications. To address this issue, we propose Ensemble Kalman Diffusion Guidance (EnKG) for diffusion models, a derivative-free approach that can solve inverse problems by only accessing forward model evaluations and a pre-trained diffusion model prior. We study the empirical effectiveness of our method across various inverse problems, including scientific settings such as inferring fluid flows and astronomical objects, which are highly non-linear inverse problems that often only permit black-box access to the forward model.
@@ 1 Introduction
The idea of using pre-trained diffusion models (Song et al., 2020; Ho et al., 2020) as plug-and-play priors for solving inverse problems has been increasingly popular and successful across various applications including medical imaging (Song et al., 2021; Sun et al., 2023), image restoration (Chung et al., 2022b; Wang et al., 2022), and music generation (Huang et al., 2024). A key advantage of this approach is its flexibility to accommodate different problems without re-training while maintaining the expressive power of diffusion models to capture complex and high-dimensional prior data distributions. However, most existing algorithms rely on privileged information of the forward model, such as its derivative (Chung et al., 2022a; Song et al., 2023b), pseudo-inverse (Song et al., 2023a), or knowledge of its parameterization (Chung et al., 2023a). This reliance poses a substantial limitation that prevents their use in problems where such information is generally unavailable. For instance, in many scientific applications (Oliver et al., 2008; Evensen & Van Leeuwen, 1996; Iglesias, 2015), the forward model consists of a system of partial differential equations whose derivative or pseudo-inverse is generally unavailable or even undefined.
@@
The goal of this work is to develop an efficient method that only requires black-box access to the forward model and pre-trained diffusion model for solving general inverse problems. Such an approach could significantly extend the range of diffusion-based inverse problems studied in the current literature, unlocking a new class of applications – especially many scientific applications. The major challenge here arises from the difficulty of approximating the gradient of a general forward model with only black-box access. The standard zero-order gradient estimation methods are known to scale poorly with the problem dimension (Berahas et al., 2022).
In this paper, we first propose a generic prediction-correction (PC) framework using an optimization perspective that includes existing diffusion guidance-based methods (Chung et al., 2022a; Song et al., 2023b; a; Peng et al., 2024; Tang et al., 2024) as special cases. The proposed PC framework consists of two main steps for each iteration: 1) the prediction step, which is a numerical integration step of the unconditional reverse ordinary differential equation (ODE) or stochastic differential equation (SDE); 2) the correction step, which is a proximal operator that maps the current point towards the high likelihood region while remaining close to the original trajectory. This viewpoint not only provides an alternative view and additional insights of the existing methods but also allows us to introduce new designs to develop a fully derivative-free guidance method. Our approach, called Ensemble Kalman Diffusion Guidance (EnKG), uses an ensemble of particles to estimate the guidance term. A key technical step is to use the empirical covariance matrix of this ensemble instead of relying on a usual scalar weight for the $L^{2}$ regularization in the correction step. Using this step allows us to derive the update rule for each ensemble member without the need for gradient, following the same approach from classic ensemble Kalman methods (Evensen, 2003; Schillings & Stuart, 2017). Consequently, EnKG is a fully derivative-free diffusion guidance method, and can solve challenging inverse problems that only permit black-box access to the forward model.
@@
  • We evaluate on various inverse problems including the standard imaging tasks and scientific problems like the Navier-Stokes equation and black-hole imaging. On the standard imaging tasks, the proposed EnKG outperforms baseline methods on the nonlinear phase retrieval by a large margin. More importantly, EnKG remains effective and efficent on the highly nonlinear inverse problem where the gradient information is genuinely inaccessible such as Navier-Stokes, while the baseline methods fail to generate reasonable predictions.
@@ 2 Background & Problem Setting
Let $G:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denote the forward model that maps the true unobserved source ${\bm{x}}$ to observations ${\bm{y}}$. We consider the following general inverse problem given by:
@@
where we only have black-box access to $G$, which we generally assume to be non-linear. $\xi$ represents the observation noise which is typically Gaussian, i.e., $\xi\sim{\mathcal{N}}(0,\Gamma)$, and ${\bm{y}}$ represents the observation which is usually sparse, i.e., $m<n$. In this paper, solving the inverse problem amounts to finding the ${\bm{x}}$ that has high posterior $p({\bm{x}}|{\bm{y}})$, which is often expressed via Bayes’s rule as $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})\cdot p({\bm{x}})$. Here $p({\bm{x}})$ is the prior distribution over source signals (which we instantiate using a pre-trained diffusion model), and $p({\bm{y}}|{\bm{x}})$ is defined as (1). Because we only have black-box access to $G$, we can only sample from $p({\bm{y}}|{\bm{x}})$, and do not know its functional form. For simplicity, we focus on finding the max a posteriori estimate: $\arg\max_{\bm{x}}p({\bm{y}}|{\bm{x}})\cdot p({\bm{x}})$.
@@
Diffusion models (Song et al., 2020; Karras et al., 2022) capture the prior $p({\bm{x}})$ implicitly using a diffusion process, which includes a forward process and backward process. The forward process transforms a data distribution ${\bm{x}}_{0}\sim p_{\mathrm{data}}$ into a Gaussian distribution ${\bm{x}}_{T}\sim{\mathcal{N}}(0,\sigma^{2}(T){\bm{I}})$ defined by a pre-determined stochastic process. The Gaussian distribution is often referred to as noise, and so the forward process ($t$ going from 0 to $T$) is typically used to create training data (iteratively noisier versions of ${\bm{x}}_{0}\sim p_{\mathrm{data}}$) for the diffusion model. The backward process ($t$ going from $T$ to 0), which is typically learned in a diffusion model, is the standard generative model and operates by sequentially denoising the noisy data into clean data, which can be done by either a probability flow ODE or a reverse-time stochastic process. In this paper, we consider the following probability flow ODE throughout the whole paper for light notation without loss of generality because every other probability flow ODE is equivalent to it up to a simple reparameterization as shown by  Karras et al. (2022):
@@
Training a diffusion model amounts to training the so-called score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$, which we assume is already completed. Given a trained diffusion model, we can sample $p({\bm{x}})$ by integrating (2) starting from a random noise sample. As such, the focus of this paper is not on training diffusion models, but rather on guiding pre-trained diffusion models for solving inverse problems, discussed next.
@@
Arguably the most popular approach to solving inverse problems with a pre-trained diffusion model is guidance-based (Chung et al., 2022a; Wang et al., 2022; Kawar et al., 2022; Song et al., 2023a; Zhu et al., 2023; Rout et al., 2023; Chung et al., 2023b; Tang et al., 2024). Guidance-based methods are originally interpreted as the conditional reverse diffusion process targeting the posterior distribution. For ease of notation and clear presentation, we use the probability flow ODE to represent the reverse process and rewrite it with Bayes Theorem.
@@
where $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ is the unconditional score and the $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$ is the guidance from likelihood. In practice, the unconditional score is approximated by a pre-trained diffusion model $s_{\theta}({\bm{x}}_{t},t)$ and the weight of the likelihood term is considered as a hyperparameter. The corresponding reverse dynamics is given by
@@
One key issue with Eq. (4) is that many algorithms for sampling along Eq. (4) assume the ability to access the gradient $\nabla_{{\bm{x}}_{t}}\log\hat{p}_{t}({\bm{y}}|{\bm{x}}_{t})$. When this gradient is unavailable (e.g., when one only has black-box access to $\hat{p}_{t}({\bm{y}})$), then one must develop a derivative-free approach, which is the central technical contribution of our work.
@@
Our approach relies on derivative-free optimization in order to solve the inverse problem Derivative-free optimization (DFO) refers to the optimization of a function where access to its derivative is not available, and only black-box access is available. Traditional DFO algorithms are broadly classified as direct search, which includes the coordinate search (Fermi, 1952), stochastic finite-difference approximations of the gradient (Spall, 1998), and Nelder-Mead simplex methods (Nelder & Mead, 1965), and model-based methods which include descent and trust region methods (Conn et al., 2000; Bortz & Kelley, 1998). Recent stochastic zero-order optimization techniques involve approximating the gradient via Gaussian smoothing (Nesterov & Spokoiny, 2017); these gradient estimates can be plugged into algorithms like stochastic gradient descent (SGD) and directly parallel their first-order counterparts (LeCun et al., 1998). Our approach builds on top of Ensemble Kalman Inversion (EKI) which is a popular scientific computing method used primarily for solving physical inverse problems (Iglesias et al., 2013; Calvello et al., 2022). From an optimization perspective, the method can be seen as performing gradient decent with a particle-based approximation to the derivative of the forward operator (Schillings & Stuart, 2017; Kovachki & Stuart, 2019).
@@ 3.1 Prediction-correction interpretation of guidance-based methods
Figure 1: Illustration of the prediction-correction interpretation for guidance-based methods on a 1D Gaussian mixture example. From left to right, the probability flow ODE gradually transforms $p_{t}({\bm{x}}_{t})$ from a Gaussian into a mixture of two Gaussians. The grey lines indicate the vector field of the probability flow. The prediction step is simply a numerical integration step over the probability flow trajectory. The correction step moves towards the MAP estimator while staying near to the initial prediction point.
@@
where the larger guidance scale $w_{i}$ gives the solution point near the MAP estimator and smaller value leads to small movement towards the MAP estimator. Eq. (6) is essentially a proximal operator (Parikh et al., 2014) if $w_{i}$ is lower bounded by a positive number. Thus it is firmly non-expansive in the convex setting and will converge to MAP estimator when applied recursively.
To solve Eq. (6) efficiently in practice, we may use first-order Taylor approximation of $\log\hat{p}({\bm{y}}|{\bm{x}}_{i+1})$ at ${\bm{x}}_{i}^{\prime}$, which gives:
$\log\hat{p}({\bm{y}}|{\bm{x}}_{i+1})=\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})+\nabla^{\top}\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})\left({\bm{x}}_{i+1}-{\bm{x}}_{i}^{\prime}\right)+O(\Delta{\bm{x}}_{i}^{2}).$ (7)
@@
${\bm{x}}_{i+1}$ $\approx\arg\min_{{\bm{x}}_{i+1}}\frac{\|{\bm{x}}_{i+1}-{\bm{x}}^{\prime}_{i}\|^{2}_{2}}{2w_{i}}-\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})-\nabla^{\top}\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})\left({\bm{x}}_{i+1}-{\bm{x}}_{i}^{\prime}\right)$ (8)
$={\bm{x}}_{i}^{\prime}+w_{i}\nabla\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime}),$ (9)
@@
Figure 1 visually demonstrates the Prediction-Correction interpretation in a 1D Gaussian mixture example. This depiction shows how the guidance-based methods iteratively step towards the MAP estimator while staying close to the initial unconditional generation trajectory defined by the prediction step. More importantly, this PC framework allows more degrees of freedom in design. In the next section, we introduce ensemble covariance matrix to the step size and derive the update rule that does not require gradient information. To stimulate further research, Section 6 discusses other potential design innovations within the PC framework.

### v2

@@ Abstract
When solving inverse problems, one increasingly popular approach is to use pre-trained diffusion models as plug-and-play priors. This framework can accommodate different forward models without re-training while preserving the generative capability of diffusion models. Despite their success in many imaging inverse problems, most existing methods rely on privileged information such as derivative, pseudo-inverse, or full knowledge about the forward model. This reliance poses a substantial limitation that restricts their use in a wide range of problems where such information is unavailable, such as in many scientific applications. We propose Ensemble Kalman Diffusion Guidance (EnKG), a derivative-free approach that can solve inverse problems by only accessing forward model evaluations and a pre-trained diffusion model prior. We study the empirical effectiveness of EnKG across various inverse problems, including scientific settings such as inferring fluid flows and astronomical objects, which are highly non-linear inverse problems that often only permit black-box access to the forward model. We open-source our code at https://github.com/devzhk/enkg-pytorch.
@@ 1 Introduction
The idea of using pre-trained diffusion models (Song et al., 2020; Ho et al., 2020) as plug-and-play priors for solving inverse problems has been increasingly popular and successful across various applications including medical imaging (Song et al., 2021; Sun et al., 2023), image restoration (Chung et al., 2022b; Wang et al., 2022), and image and music generation (Rout et al., 2024b; Huang et al., 2024). A key advantage of this approach is its flexibility to accommodate different problems without re-training while maintaining the expressive power of diffusion models to capture complex and high-dimensional prior data distributions. However, most existing algorithms rely on privileged information of the forward model, such as its derivative (Chung et al., 2022a; Song et al., 2023b), pseudo-inverse (Song et al., 2023a), or knowledge of its parameterization (Chung et al., 2023a). This reliance poses a substantial limitation that prevents their use in problems where such information is generally unavailable. For instance, in many scientific applications (Oliver et al., 2008; Evensen & Van Leeuwen, 1996; Iglesias, 2015), the forward model consists of a system of partial differential equations whose derivative or pseudo-inverse is generally unavailable or even undefined.
@@
The goal of this work is to develop an effective method that only requires black-box access to the forward model and pre-trained diffusion model for solving general inverse problems. Such an approach could significantly extend the range of diffusion-based inverse problems studied in the current literature, unlocking a new class of applications – especially many scientific applications. The major challenge here arises from the difficulty of approximating the gradient of a general forward model with only black-box access. The standard zero-order gradient estimation methods are known to scale poorly with the problem dimension (Berahas et al., 2022).
To develop our approach, we first propose a generic prediction-correction (PC) framework using an optimization perspective that includes existing diffusion guidance-based methods (Chung et al., 2022a; Song et al., 2023b; a; Peng et al., 2024; Tang et al., 2024) as special cases. The key idea of this PC framework is to decompose diffusion guidance into two explicitly separate steps, unconditional generation (i.e., sampling from the diffusion model prior), and guidance imposed by the observations and forward model. This modular viewpoint allows us to both develop new insights of the existing methods, as well as to introduce new tools to develop a fully derivative-free guidance method. Our approach, called Ensemble Kalman Diffusion Guidance (EnKG), uses an ensemble of particles to estimate the guidance term while only using black-box queries to the forward model (i.e., no derivatives are needed), using a technique known as statistical linearization (Evensen, 2003; Schillings & Stuart, 2017) that we introduce to diffusion guidance via our PC framework.
@@
  • We evaluate on various inverse problems including the standard imaging tasks and scientific problems like the Navier-Stokes equation and black-hole imaging. On more challenging tasks, such as nonlinear phase retrieval in standard imaging and the scientific inverse tasks, our proposed EnKG outperforms baseline methods by a large margin. For problems with very expensive forward models (e.g., Navier-Stokes equation), EnKG also stands out as being much more computationally efficient than other derivative-free methods.
@@ 2 Background & Problem Setting
Let $G:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denote the forward model that maps the true unobserved source ${\bm{x}}$ to observations ${\bm{y}}$. We consider the following setting:
@@
where we only have black-box access to $G$ (generally assumed to be non-linear). $\xi$ represents the observation noise which is often modeled as Gaussian, i.e., $\xi\sim{\mathcal{N}}(0,\Gamma)$, and ${\bm{y}}$ represents the observation. Solving the inverse problem amounts to inverting Eq. (1), i.e., finding the most likely ${\bm{x}}$ (MAP inference) or its posterior distribution $P({\bm{x}}|{\bm{y}})$ (posterior inference) given ${\bm{y}}$. This inverse task is often expressed via Bayes’s rule as $p({\bm{x}}|{\bm{y}})\propto p({\bm{y}}|{\bm{x}})\cdot p({\bm{x}})$. Here $p({\bm{x}})$ is the prior distribution over source signals (which we instantiate using a pre-trained diffusion model), and $p({\bm{y}}|{\bm{x}})$ is defined as (1). Because we only have black-box access to $G$, we can only sample from $p({\bm{y}}|{\bm{x}})$, and do not know its functional form. For simplicity, we focus on finding the MAP estimate: $\arg\max_{\bm{x}}p({\bm{y}}|{\bm{x}})\cdot p({\bm{x}})$.
@@
Diffusion models (Song et al., 2020; Karras et al., 2022) capture the prior $p({\bm{x}})$ implicitly using a diffusion process, which includes a forward process and backward process. The forward process transforms a data distribution ${\bm{x}}_{0}\sim p_{\mathrm{data}}$ into a Gaussian distribution ${\bm{x}}_{T}\sim{\mathcal{N}}(0,\sigma^{2}(T){\bm{I}})$ defined by a pre-determined stochastic process. The Gaussian distribution is often referred to as noise, and so the forward process ($t$ going from 0 to $T$) is typically used to create training data (iteratively noisier versions of ${\bm{x}}_{0}\sim p_{\mathrm{data}}$) for the diffusion model. The backward process ($t$ going from $T$ to 0), which is typically learned in a diffusion model, is the standard generative model and operates by sequentially denoising the noisy data into clean data, which can be done by either a probability flow ODE or a reverse-time stochastic process. Without loss of generality, we consider the following probability flow ODE since every other probability flow ODE is equivalent to it up to a simple reparameterization as shown by  Karras et al. (2022):
@@
Training a diffusion model amounts to training the so-called score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$, which we assume is already completed (and not the focus of this paper). Given a trained diffusion model, we can sample $p({\bm{x}})$ by integrating (2) starting from random noise.
@@
As surveyed in Daras et al. (2024), arguably the most popular approach to solving inverse problems with a pre-trained diffusion model is guidance-based (Chung et al., 2022a; Wang et al., 2022; Kawar et al., 2022; Song et al., 2023a; Zhu et al., 2023; Rout et al., 2023; Chung et al., 2023b; Tang et al., 2024). Guidance-based methods are originally interpreted as the conditional reverse diffusion process targeting the posterior distribution. For ease of notation and clear presentation, we use the probability flow ODE to represent the reverse process and rewrite it with Bayes Theorem.
@@
where $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ is the unconditional score and the $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{y}}|{\bm{x}}_{t})$ is the guidance from likelihood. In practice, the unconditional score is approximated by a pre-trained diffusion model $s_{\theta}({\bm{x}}_{t},t)$. The likelihood term, $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int_{{\bm{x}}_{0}}p({\bm{x}}_{t}|{\bm{x}}_{0})p({\bm{y}}|{\bm{x}}_{0}){\mathrm{d}}{\bm{x}}_{0}$, is computationally intractable as it requires integration over all possible ${\bm{x}}_{0}$. Various tractable approximations have been proposed in the literature, which we denote as $\hat{p}_{t}({\bm{y}}|{\bm{x}}_{t})$. The corresponding approximated reverse process is:
@@
One key issue with Eq. (4) is that many algorithms for sampling along Eq. (4) assume access to the gradient $\nabla_{{\bm{x}}_{t}}\log\hat{p}_{t}({\bm{y}}|{\bm{x}}_{t})$. When this gradient is unavailable (e.g., only black-box access to $\hat{p}_{t}({\bm{y}})$), then one must develop a derivative-free approach, which is our core technical contribution.
Two existing derivative-free diffusion guidance methods are stochastic control guidance (SCG) (Huang et al., 2024), and diffusion policy gradient (DPG) (Tang et al., 2024). Both SCG and DPG are developed from the stochastic control viewpoint, and guides the diffusion process via estimating a value function, which can be challenging to estimate well (as seen in our experiments).
3 Related work
Ensemble Kalman methods
Ensemble Kalman methodology was first introduced by Evensen (1994) in the context of data assimilation and later revisited by Iglesias et al. (2013) for inverse problems from an optimization perspective, resulting in the derivative-free algorithm known as Ensemble Kalman Inversion (EKI). Subsequent advancements include momentum-based EKI for training neural networks (Kovachki & Stuart, 2019) and various regularization techniques to improve stability and efficiency (Iglesias, 2016; Chada et al., 2020). While the original derivations of these methods (Evensen, 1994; Iglesias et al., 2013) may not explicitly highlight it, the overarching idea behind the ensemble Kalman methodology can be interpreted as linearization of the forward model $G$ as elucidated in Schillings & Stuart (2017); Law et al. (2016). In essence, we approximate the potentially complex forward model $G(x)$ with the best-fitting linear surrogate model $y=Ax+b$. In other words, we seek the minimizers of the following optimization problem:
$\min_{A,b}\mathbb{E}_{x}\|G(x)-Ax-b\|_{2}^{2},$
which has closed-form solutions given by
$A=\mathbb{E}_{x}[(G(x)-\mathbb{E}_{x}G(x))x^{\top}]C_{xx}^{-1},b=\mathbb{E}_{x}[G(x)]-\mathbb{E}_{x}[Ax],$
where $C_{xx}^{-1}$ is the pseudoinverse of the covariance matrix. Our approach builds upon this linear surrogate model to approximate the forward model in our likelihood step defined in Eq. (14), enabling a fully derivative-free algorithm. Furthermore, prior works (Bergou et al., 2019; Chada & Tong, 2022) establish the convergence results for certain EKI variants in the non-linear setting. However, these proofs do not directly apply to our algorithm. Instead, we develop tailored convergence results for our analysis.
@@ 3 Related work
Traditional derivative-free optimization (DFO) algorithms include direct search, which includes the coordinate search, stochastic finite-difference approximations of the gradient, Nelder-Mead simplex methods, and model-based methods; see Berahas et al. (2022) for an overview. Among modern DFO techniques, Gaussian smoothing methods (Nesterov & Spokoiny, 2017) have demonstrated robust empirical performance (Salimans et al., 2017). These gradient estimates can be plugged into gradient-based algorithms directly, which we use to establish strong baselines in this paper.
Predictor-Corrector method for diffusion models
The term "predictor-corrector" has been used in several prior works on diffusion model sampling, but their objectives and mechanisms differ from our PC framework. Song et al. (2020) introduces a Predictor-Corrector sampler for the sampling of diffusion models, where both the predictor and corrector aim to sample from the same target distribution by simulating different stochastic processes (reverse-time SDE and annealed Langevin dynamics, respectively). Lezama et al. (2022) then extends this framework to discrete space. More recently, Bradley & Nakkiran (2024) applies a similar perspective to classifier-free guidance, using PF-ODE solver (DDIM) as the predictor and Langevin dynamics as corrector, in order to sample from the gamma-powered data distribution. Other works, such as Zhao et al. (2023) and Zhao et al. (2024), draw inspiration from predictor-corrector methods in the literature of classical numerical ODE solvers, focusing on solving the probability flow ODE (PF-ODE) with higher-order accuracy and adaptive step sizes. These existing PC methods aim to sample from the diffusion model, which is related to the prediction step (sampling from diffusion prior) in our framework but irrelevant to our correction step that interacts with the forward model $G$.
@@ 4.1 Prediction-correction interpretation of guidance-based methods
Prior prediction step
Figure 1: Illustration of the prediction-correction interpretation for guidance-based methods on a 1D Gaussian mixture example. From left to right, the probability flow ODE transforms $p_{t}({\bm{x}}_{t})$ from a Gaussian into a mixture of two Gaussians. The grey lines represent the trajectories of the probability flow. The prediction step corresponds to a numerical integration step along this flow. The correction step adjusts the particle towards the MAP estimator while keeping them close to the initial prediction point. In contrast to prior methods with independent particles, the proposed EnKG introduces interactions between particles to eliminate the need for gradient access.
@@
where the larger guidance scale $w_{i}$ gives the solution point near the MAP estimator and smaller value leads to small movement towards the MAP estimator. Eq. (6) is essentially a proximal operator (Parikh et al., 2014) if $w_{i}$ is lower bounded by a positive number. This optimization problem is often non-convex in most practical scenarios. As a result, the optimization algorithm may converge to a local maximum rather than a global one.
To solve Eq. (6) efficiently, one can use a first-order Taylor approximation of $\log\hat{p}({\bm{y}}|{\bm{x}}_{i+1})$ at ${\bm{x}}_{i}^{\prime}$:
$\log\hat{p}({\bm{y}}|{\bm{x}}_{i+1})=\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})+\nabla_{{\bm{x}}}^{\top}\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})\left({\bm{x}}_{i+1}-{\bm{x}}_{i}^{\prime}\right)+O\left(|{\bm{x}}_{i+1}-{\bm{x}}_{i}|^{2}\right).$ (7)
@@
${\bm{x}}_{i+1}$ $\approx\arg\min_{{\bm{x}}_{i+1}}\frac{\|{\bm{x}}_{i+1}-{\bm{x}}^{\prime}_{i}\|^{2}_{2}}{2w_{i}}-\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})-\nabla_{{\bm{x}}}^{\top}\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime})\left({\bm{x}}_{i+1}-{\bm{x}}_{i}^{\prime}\right)$ (8)
$={\bm{x}}_{i}^{\prime}+w_{i}\nabla_{{\bm{x}}}\log\hat{p}({\bm{y}}|{\bm{x}}_{i}^{\prime}),$ (9)
@@
Figure 1 depicts the Prediction-Correction interpretation in a 1D Gaussian mixture example, where guidance-based methods iteratively step towards the MAP estimator while staying close to the initial unconditional generation trajectory defined by the prediction step. Importantly, the PC framework allows more degrees of freedom in method design.